# Motion-Aware Animatable Gaussian Avatars Deblurring

### v2

@@ Abstract
The creation of 3D human avatars from multi-view videos is a significant yet challenging task in computer vision. However, existing techniques rely on high-quality, sharp images as input, which are often impractical to obtain in real-world scenarios due to variations in human motion speed and intensity. This paper introduces a novel method for directly reconstructing sharp 3D human Gaussian avatars from blurry videos. The proposed approach incorporates a 3D-aware, physics-based model of blur formation caused by human motion, together with a 3D human motion model designed to resolve ambiguities in motion-induced blur. This framework enables the joint optimization of the avatar representation and motion parameters from a coarse initialization. Comprehensive benchmarks are established using both a synthetic dataset and a real-world dataset captured with a 360-degree synchronous hybrid-exposure camera system. Extensive evaluations demonstrate the effectiveness and robustness of the model across diverse conditions.
@@ 1 Introduction
In the field of 3D reconstruction, creating high-quality 3D human avatars holds immense potential for the industry [19, 68, 14, 35, 4, 15, 39, 84, 75, 5, 50, 81]. Cutting-edge methods employing 3D Gaussian Splatting [22, 9, 41, 10] and the Skinned Multi-Person Linear (SMPL) model [42] have achieved notable success. These techniques [14, 57, 13, 35] typically rely on video data with SMPL parameters from calibration techniques like EasyMocap [7, 56].
@@
While these models have achieved remarkable results with well-calibrated and high-quality video frames, their performance can be significantly degraded by blur effects caused by human motion. In practical scenarios, such motion-induced blur is often unavoidable due to the unpredictable speeds and movements of subjects. In particular, blur effects can adversely affect the performance of existing models in two key ways. First, blurry captures may cause the 3DGS model to learn a distorted 3D representation, arising from the inherent ambiguity of motion blur (Fig. 1). Such ambiguity hinders the accurate recovery of structural information and texture details. Second, even though static cameras can be calibrated before recording, blurry captures still lead to erroneous estimation of SMPL parameters. [^1] Some studies, such as [14], incorporate a correction module to refine pose parameters during training. However, experimental results indicate that these modifications often lead to suboptimal performance due to the absence of a precise blur-aware motion modeling process.
@@ 3 Method
Sub-frame rigid sequential pose model. The pose parameters $\bm{\Theta}_{t}$, defined in the SMPL model as $\bm{\Theta}_{t}\in\mathbb{R}^{24\times 3}$, represent the rotations of $24$ joints at each time step $t$, expressed in $\textbf{SO}(3)$. Considering the inherent continuity of joint motion, De Boor–Cox formulation for B-splines [58, 8, 66, 33] is employed to interpolate intermediate poses.
@@ 4 Experiments
4.1 Benchmarks
Synthetic dataset. The synthetic data is constructed based on the ZJU-MoCap [56] dataset since it is widely adopted by many 3D human avatar research work [14, 11, 72, 57, 23, 77]. Following their practice, six human subjects (IDs: $377$, $386$, $387$, $392$, $393$, $394$) are selected for experiments. Each blurry frame was synthesized by averaging $K_{blur}$ consecutive sharp frames To mitigate discontinuities caused by direct averaging, the RIFE model [16] is employed to interpolate $16$ intermediate frames between adjacent sharp frames. Blurry frames from Cameras $04$, $10$, $16$, and $22$ are used for training, while the sharp frames from the remaining $18$ cameras are reserved for novel-view test. SMPL parameters are calibrated using the blurry video with EasyMoCap [7, 56], and human masks are obtained using SAM [25].
@@ 4.1 Benchmarks
Real dataset. The dataset consists of real-captured blurry videos and synchronously captured sharp videos from novel viewpoints. The capture system employs $12$ time-synchronized Blackfly BFS-U3-51S5C cameras arranged in a 360-degree spatial configuration, as illustrated in Fig. 3. $4$ cameras record blurry videos, while the remaining $8$ capture sharp videos for novel-view evaluation. Fig. 5 illustrates the time synchronization of two types of cameras. To ensure consistent pixel intensity, the blurry cameras are equipped with a 1/16{1}/{16} ND filter. Detailed configuration are reported in Tab. 1. SMPL parameters are calibrated using EasyMoCap [7], and human masks are generated using SAM [25].

### v3

@@ Abstract
The creation of 3D human avatars from multi-view videos is a significant yet challenging task in computer vision. However, existing techniques rely on high-quality, sharp images as input, which are often impractical to obtain in real-world scenarios due to variations in human motion speed and intensity. This paper introduces a novel method for directly reconstructing sharp 3D human Gaussian avatars from blurry videos. The proposed approach incorporates a 3D-aware, physics-based model of blur formation caused by human motion, together with a 3D human motion model designed to resolve ambiguities in motion-induced blur. This framework enables the joint optimization of the avatar representation and motion parameters from a coarse initialization. Comprehensive benchmarks are established using both a synthetic dataset and a real-world dataset captured with a 360-degree synchronous hybrid-exposure camera system. Extensive evaluations demonstrate the effectiveness of the model across diverse conditions. Codes Available: https://github.com/MyNiuuu/MAD-Avatar
@@ 1 Introduction
In the field of 3D reconstruction, creating high-quality 3D human avatars holds immense potential for the industry [18, 68, 13, 34, 4, 14, 38, 84, 75, 5, 49, 81]. Cutting-edge methods employing 3D Gaussian Splatting [21, 8, 40, 9] and the Skinned Multi-Person Linear (SMPL) model [41] have achieved notable success. These techniques [13, 57, 12, 34] typically rely on video data with SMPL parameters from calibration techniques like EasyMocap [56].
@@
While these models have achieved remarkable results with well-calibrated and high-quality video frames, their performance can be significantly degraded by blur effects caused by human motion. In practical scenarios, such motion-induced blur is often unavoidable due to the unpredictable speeds and movements of subjects. In particular, blur effects can adversely affect the performance of existing models in two key ways. First, blurry captures may cause the 3DGS model to learn a distorted 3D representation, arising from the inherent ambiguity of motion blur (Fig. 1). Such ambiguity hinders the accurate recovery of structural information and texture details. Second, even though static cameras can be calibrated before recording, blurry captures still lead to erroneous estimation of SMPL parameters. [^1] While certain approaches, such as GauHuman [13], integrate pose refinement modules, they often yield suboptimal results because they lack a rigorous mechanism for modeling blur-aware motion.
@@ 3 Method
Sub-frame rigid sequential pose model. The pose parameters $\bm{\Theta}_{t}$, defined in the SMPL model as $\bm{\Theta}_{t}\in\mathbb{R}^{24\times 3}$, represent the rotations of $24$ joints at each time step $t$, expressed in $\textbf{SO}(3)$. Considering the inherent continuity of joint motion, De Boor–Cox formulation for B-splines [58, 7, 66, 32, 50] is employed to interpolate intermediate poses.
@@ 4 Experiments
4.1 Benchmarks
Synthetic dataset. The synthetic data is constructed based on the ZJU-MoCap [56] dataset since it is widely adopted by many 3D human avatar research work [13, 10, 72, 57, 22, 77]. Following their practice, six human subjects (IDs: $377$, $386$, $387$, $392$, $393$, $394$) are selected for experiments. Each blurry frame was synthesized by averaging $K_{blur}$ consecutive sharp frames To mitigate discontinuities caused by direct averaging, the RIFE model [15] is employed to interpolate $16$ intermediate frames between adjacent sharp frames. Blurry frames from Cameras $04$, $10$, $16$, and $22$ are used for training, while the sharp frames from the remaining $18$ cameras are reserved for novel-view test. SMPL parameters are calibrated using the blurry video with EasyMoCap [56], and human masks are obtained using SAM [24].
@@ 4.1 Benchmarks
Real dataset. The dataset consists of real-captured blurry videos and synchronously captured sharp videos from novel viewpoints. The capture system employs $12$ time-synchronized Blackfly BFS-U3-51S5C cameras arranged in a 360-degree spatial configuration, as illustrated in Fig. 3. $4$ cameras record blurry videos, while the remaining $8$ capture sharp videos for novel-view evaluation. Fig. 5 illustrates the time synchronization of two types of cameras. To ensure consistent pixel intensity, the blurry cameras are equipped with a 1/16{1}/{16} ND filter. Detailed configuration are reported in Tab. 1. SMPL parameters are calibrated using EasyMoCap, and human masks are generated using SAM [24].
@@ 5 Conclusion
Acknowledgment
This research was supported in part by the Tateisi Science and Technology Foundation, JSPS KAKENHI Grant Numbers 24KK0209 and 22H00529, and Advanced AI Talent Development to Lead the Next-Generation AI for Intelligent Society (BOOST NAIS) of The University of Tokyo.